verify that each given function is a solution of the

differential equation.

7.ty′\:−\:y\:=\:t^2;\:y\:=\:3t\:+\:t^2


1
Expert's answer
2021-09-02T16:39:58-0400

Let us verify that the function "y(t)=3t+t^2" is a solution of the differential equation "ty'-y=t^2". taking into account that "y'(t)=3+2t" and "ty'(t)-y(t)=t(3+2t)-(3t+t^2)=3t+2t^2-3t-t^2=t^2," we conclude that the function "y(t)=3t+t^2" is a solution of the differential equation "ty'-y=t^2".